Version 7 (modified by mhuber@…, 4 months ago)

--

PS1 IPP Czar Logs for the week YYYY.MM.DD - YYYY.MM.DD

(Up to PS1 IPP Czar Logs)

Monday : 2017.08.14

  • MEH: Haydn fixed the i03-con issue so ipp105-112 are accessible again -- setting ipp105 neb-host up again

Tuesday : YYYY.MM.DD

Wednesday : YYYY.MM.DD

Thursday : 2017.08.17

  • MEH: registratiopn backed up from o7982g0094o not clearing summitcopy -- stalled at 60 imfiles, so did revertcopied and cleared -- things are moving along now
  • MEH: QUB targeted nightly processing running ~1430-1630 this afternoon in nightly pantasks after the backup last night --

Friday : YYYY.MM.DD

Saturday : YYYY.MM.DD

Sunday : 2017.08.20

  • 10:40 EAM : 5 end-of-night darks failed to download. for 3 exposures, there were failures in pzDownloadImfile (fault = 202), while for 2 exposures the faults were in summitExp (still 202). I cleared these manually in mysql (should add pztool options to avoid mistakes).
  • 10:50 EAM : looks like we are still having trouble reading the file listings for two exposures (c7985g0081d and c7985g0082d). conductor hangs when trying to download the index for either of these exposures. Sent a message requesting help to ps-camera.
  • MEH: QUB updates will be running during day through early part of week (~ippqub/stdscience/ptolemy.rc, ~ippqub/stdscience_ws/ptolemy.rc)